OpenStudy (anonymous):

Write the equation of the line in slope-intercept form that passes through the point (3, -2) and has a slope of 2. A. y = 2x - 8 B. y = 2x - 4 C. y = 2x + 7 D.y = 2x + 5

OpenStudy (anonymous):

y-y1 = m(x-x1) here y1 = -2 x1 = 3 and m = 2 do it now

OpenStudy (ryan123345):

u need help?

OpenStudy (anonymous):

Ok but where di you het "m" out from?

OpenStudy (anonymous):

*did * get and yeah kinda

OpenStudy (anonymous):

m is the slope, which in this case is 2. y-y1=m(x-x1) M=2 y1=-2 x1=3